Be sure to answer all parts. The key step in the metabolism of glucose for energy is the isomerization of glucose-6-phosphate (G6P) to fructose-6-phosphate (F6P): G6P_F6P K = 0.510, at 298K. Calculate AG when Q, the [F6P|/[G6P] ratio, equals 10.0. x10 J/mol Enter your answer in scientific notation.
Answer #1

dG0 = -RT ln K

Putting values:

dG0 = -8.314*298*ln(0.510) = 1668.26 J/mol

Using relation:

dG = dG0 + RT ln Q

Putting values:

dG = 1668.26 + 8.314*298*ln(10) = 7373.08 J = 7.373*103 J/mol